OpenStudy (anonymous):

Find the general solution of the homogeneous differential equation (x^2-y^2)dx+(3xy)dy=0

OpenStudy (anonymous):

\[y=\pm \sqrt{2c1x ^{2/3}-x ^{2}}/\sqrt{2}\]

OpenStudy (anonymous):

could i write that as: (x^2+2y^2)^3=cx^2

OpenStudy (anonymous):

??

OpenStudy (anonymous):

this is not a exact equation

OpenStudy (anonymous):

sorry, should have been y(x)=

OpenStudy (anonymous):

I think u sub y= vx, it can be made manageable.

OpenStudy (anonymous):

U will get a separable equation.

OpenStudy (anonymous):

->(x^2-vx^2)dx +3vx^2 dy (and dy= vdx+xdv) then rearrange.